Cos^2 π/12 +cos^2 π/4 +cos^2 5π/12 is:. ??????????????????????!???????

Grade:11

1 Answers

Venky
7 years ago
cos^2 15+cos ^2 45+cos ^2 75 =((root 6 +root2)/4)^2 +(1/root2)^2+((root6-root2)/4)^2         =(6+2+2root12)/16 +1/2 +(6+2-2root12)/16=1+(1/2)=3/2